Instrument Log 

BLC†††††††††††††††† C1††††††††††††††††† Lamont, OK Optical Trailer

Date

Operator

Comments

5/23/00

MR

The BLC has been turned off and the transmitter board removed. Belfort has indicated that they can probably repair the board. The mentor is waiting for an RMA number from Belfort. Once we have the RMA, the transmitter board will need to be packed and shipped to Belfort. I'll get the address with the RMA.

5/3/00

KT

At 1850 GMT observed that the laptop was not displaying any graphical data. At 1854 GMT powered the instrument off. At 1855 GMT powered the instrument on. RTS at 1907 GMT.

4/18/00

MR

No data was being collected from the instrument. T/S indicated the AC power was out. Checked the UPS and found it dead. Power cycled the UPS and the AC power returned. System came up and started to take data at 1350.

1/6/2000

CM

While retiling the Optical trailer floor, contractors plugged a compressor into a UPS outlet which caused the breaker to trip. The UPS and all Optical trailer instruments lost power and had to be restarted. Removed the compressor from the UPS outlet and reset the breaker. Restarted all Optical trailer instruments. RTS status O.K.

12/28/99

MK

Due to power line maintenance starting at 1500 GMT all instruments will be down unless battery power is present. At 1536 GMT the AC power was restored to all CF instruments. The BLC laptop was rebooted at 1552 GMT. RTS at 1552 GMT.

12/20/99

KT

The BLC was not collecting data. At 2155 GMT power cycled the laptop as well as the instrument. RTS at 2205 GMT.

11/1/99

CAM

The BLC laptop computer time was one hour off following the change from daylight savings time. Set the BLC laptop time at 1446 GMT. The time was also incorrect in the Clouds software. Exited the Clouds program and entered terminal mode at 2002. Set the time in the Clouds program to 2002 GMT. Exited terminal mode and program RTS at 2003 GMT.

10/4/99

GRS

SDS informed that the BLC had not sent any files since the 3rd. Needed to troubleshoot. Troubleshooting found that the BLC program had locked up on Sunday the 3rd at 1227 GMT. Stopped the Clouds program at 1810 GMT and then restarted it. This reset the software but did not correct the problem. Shut the computer down and rebooted the computer at 1813 GMT. This did not fix the problem. Powered off the instrument at 1823 GMT and then back on at 1824 GMT. This fixed the problem. RTS system at 1825

7/20/99

MW

The BLC laptop computer time was about 4 minutes, 30 seconds faster than actual time. Reset the BLC laptop time at 1404 GMT.

6/2/99

MW

The BLC system time was fast by more that four minutes. Reset system time at 1601 GMT. RTS at 1602 GMT.

4/22/99

MR

Rebooted at 1500 due to power outage.

4/1/99

MW

The BLC laptop display looked abnormal and there was no signal displayed. Restarted the laptop computer at 1415 GMT. The display returned to normal after the restart. RTS at 1418 GMT.

3/16/99

MW

The CF lost power due to a snow storm on 3/14/99 at 0030 GMT. The BLC has been down since that time. Powered the BLC laptop on this morning at 1445 GMT. The system is up and running as of 1450 GMT.

3/2/99

MW

The BLC laptop computer time was 3 minutes fast. Pressed (alt-T) and went into terminal mode. Reset the time at 1414 GMT. Pressed (alt-X) to exit from terminal mode. No signal appeared on the laptop display after setting the time. Had to power cycle the ceilometer at 1435 GMT. The normal signal returned after the power cycle. RTS at 1435 GMT.

2/19/99

SS

PM. Cleaned windows at 1551 GMT. Status was normal at this time.

12/28/98

MW

The BLC laptop display was blank this morning. The system appeared to be in a suspended mode because the normal display came back after pressing the power button at 1455 GMT. However, the program had been stopped at 2248 GMT on 12/25/98 according to the display. No data was being collected. Shut down all programs and re-started the laptop at 1456 GMT. Data collection resumed at approximately 1500 GMT.

12/16/98

MW

The BLC laptop time was fast by 1 minute 30 seconds. Reset laptop time at 1450 GMT.

12/1/98

CAM

The BLC laptop hard drive had completely filled up. With less than 10 Mb total disk space remaining, there was no place left to move the data files to. FTP'd all data files from e:\data to storage on another PC from 1945 GMT to 2045 GMT. Deleted all files from the e:\data directory. Moved all recent data files from c:\clouds\data to e:\data. This freed up 289 Mb on c:\clouds\data. RTS status O.K. at 2105 GMT. Notified mentor where the data was relocated.

11/24/98

MW

The BLC program clock was 5 minutes fast. Reset the terminal clock at 1437 GMT.

11/18/98

MW

The BLC laptop hard disk (C:) was full and data could not be written to the disk. The error message 'cannot write to disk' would pop up whenever the BLC sent a data packet. Moved several data files off of the C: drive onto the compressed drive to free up some space until the instrument mentor can pull the data off of the laptop. Data transfer completed at 1500 GMT. Called the instrument mentor and left a message.

10/1/98

CAM

The BLC laptop display was garbled and indicated an error message 'Unable to write data to disk' each time a data packet was received. Checked C: drive and found that it was full. Checked c:\clouds\data directory and found it contained 198 Mb of data in files from 8/26/98 to 10/01/98. Created a new 'Data' directory on E: drive. Moved all files except for today's from c:\clouds\data to e:\data from 1845 GMT to 1858 GMT. The Ceilometer clock was 5 minutes fast. Set Ceilometer clock to SDS time at 1901 GMT. Restarted Clouds program and RTS status O.K. at 1903 GMT.

9/30/98

KT

Unable to view data due to laptop being shut off. At 1605 GMT powered laptop back on and was able to view data scrolling.

9/10/98

SS

The laptop was locked up and will not respond to keyboard or mouse controls. Power cycled the computer and it come up and began updating. RTS at 1600 GMT.

7/22/98

CAM

Mentor requested that the right angle blind be placed over the BLC transmit and receive windows to further T/S the cloud base offset problem. Mentor requested this be done in stages, first fully covered then gradually uncovered in one inch increments for about 5 minutes each step. Placed the blind fully over the windows from 1825 to 1834 GMT. Moved the blind to expose 1 inch of the windows from 1836 to 1840 GMT, 2 inches from 1843 to 1848, 3 inches from 1850 to 1855, 4 inches from 1858 to 1903 and finally 5 inches exposed from 1906 to 1911 GMT. The blind was removed completely and BLC RTS at 1911 GMT.

7/22/98

MW

The BLC laptop computer display was not updating this morning. The display indicated the last packet received was at 2113 GMT on 7/21/98. Rebooted the computer at 1345 GMT. No packets received after reboot. Power cycled the BLC at 1355 GMT. Packets received after power cycle. RTS at 1355 GMT.

7/16/98

GRS

The Mentor requested that a "right angle blind" be built to bounce the laser from one telescope into the other. The laser exits one telescope, bounces off a 45 degree white target to another 45 degree white target, and then back into the other telescope. Constructed the "right angle blind" according to Conner Flynnsí specifications. Started the placement and leveling of the blind over the BLC at 1935 GMT. Had the blind in place and started the 15 minute reading from 1945 to 2000 GMT. Just for precautionary sakes, moved the blind over the telescopes one more time, leveled and proceeded with a ten minute reading from 2011 to 2021 GMT. Removed the blind at 2021 GMT. RTS BLC at 2022 GMT.

7/1/98

MW

Moisture has built up inside the instrument shelter windows. No CM performed at this time due to current precipitation.

6/22/98

MW

The BLC laptop was not updating and no data was displayed. The date on the computer read 6/20/98 at 1909 GMT. Rebooted the laptop computer at 1426 GMT. The laptop still didn't receive any data from the BLC. Power cycled the BLC at 1335 GMT. Data collection resumed after the BLC power cycle. RTS at 1335 GMT.

6/17/98

SS

Instrument mentor requested that the instrument and laptop be power cycled due to loss of network connection. At 1910 shut down laptop software and power cycled. At 1920 checked bit drivers to instrument no trouble found. At 1923 power cycled instrument. RTS at 1930 GMT.

6/16/98

SS

Work order 2894 states that the SDS reports no files to be ingested. Power cycled the instrument at 1902 GMT. RTS at 1915 GMT.

5/29/98

CAM

Instrument mentor requested that the VLC be removed from service. This required a brief power interruption to BLC as they are both on the same AC circuit. Secured BLC/VLC AC power at 2014 GMT. Disconnected VLC from power. Power to BLC restored at 2015 GMT.

5/26/98

MK

Due to a power failure on 5/25/98 at 0425 GMT, the BLC computer has shut off. Powered on the BLC computer at 1615 GMT, status OK. RTS at 1630 GMT.

5/15/98

CAM

The site lost all power due to a severe thunderstorm at approximately 1030 GMT. Site power was restored at 1400 GMT. BLC status check at 1605 GMT found that the laptop was powered off. Powered the laptop on at 1605 GMT. The last data displayed according to the time line was at 1055 GMT. Display began updating with the correct date and time at 1606 GMT. RTS status O.K.

5/7/98

CAM

The CF experienced a power outage due to a county transformer going down around 2130 GMT on 5/6/98. BLC status check on 5/7/98 found the laptop in sleep mode. Activated the laptop at 1417 GMT on 4/7 and the clouds display immediately came up. However the date and time indicated was 5/6/98, 2243 GMT. After a couple of seconds the program updated with the current date and time and the system appeared to be fully operational. A check of the SDS status page on the WEB at 1700 GMT showed the ingest process was still red with the last successful ingest being on 5/6/98 at 2251 GMT. Checked the raw data files on SDS and found that the BLC is in fact writing files. System is operational but it is unclear why SDS is not ingesting the files at this time. Will continue to monitor the status page.

5/5/98

CAM

Daily status check found that the laptop was not powered on. Powered laptop on at 1300 GMT. The program was not updating and the time showed 0947, 5/5/98. The graph updated itself and resumed normal operation at 1301 GMT. RTS status O.K. at 1305 GMT. The time of outage corresponds with a county wide power outage experienced at the CF around 0918 GMT on 5/5 that lasted approximately 2 hours.

4/28/98

MW

The BLC laptop was locked up this morning and would not allow any keyboard or mouse commands. The clock display read 1104 GMT as the time it locked up. Rebooted the laptop at 1407 GMT. The computer came back online and started receiving packets from the BLC at 1411 GMT. RTS.

4/17/98

CAM

Secured AC power to Belfort Ceilometer at 1553 GMT for installation of 25K Ceilometer. Installed 25K Ceilometer. Restored power and Belfort Ceilometer RTS at 1600 GMT.

3/9/98

SS

Found the laptop locked up. Laptop would not respond to mouse or keyboard. Cycled power. laptop display now updating. RTS at 1845 GMT.

3/6/98

SS

Found the laptop not updating or responding to the keyboard or mouse. Power cycled the laptop. Now the display is normal. RTS at 1715 GMT.

3/5/98

SS

Found the laptop not updating or responding to the mouse or keyboard. Cycled laptop power. Laptop is now updating and responding correctly. RTS at 1745 GMT.

2/19/98

KT

Laptop was not updating the data. At 1640 GMT the laptop was power cycled. At 1657 GMT the BLC was power cycled. Laptop came back up and data started scrolling. RTS at 1700 GMT.

1/23/98

SS

Found laptop with no graph displayed, zeros for date and time. Performed a SW reboot on laptop and restarted program with the same results at 1625. Power cycled the Ceilometer at 1635. Reset the time at 1642. RTS at 1642.

1/20/98

MW

The BLC was reporting the date as 1/21/98 instead of 1/20/98. The laptop computer time and date was correct. Typed 'alt-T' on the laptop and went into terminal mode at 1750 GMT. Changed the date to 1/20/98. Returned to 'Run-mode' at 1752 GMT. Date is now correct. RTS.

1/14/98

KT

The BLC laptop was not updating the data nor was it showing the cloud heights. At 1818 GMT the laptop was re-booted. Still not operating. At 1842 GMT power cycled the system. Data started to update on the laptop and the cloud heights were available. The cloud height reported at 1125 ft, which is agreeable with the outside conditions. RTS at 1845 GMT.

1/13/98

MW

The BLC was not communicating with the laptop computer. No graph was displayed on the laptop and the clock was stopped at 1550 GMT on 1/9/98. Disconnected the BLC bit driver from the BLC RS-232 connector and placed a plug with pins 2 and 3 shorted together on the bit driver. Performed a loop back test between the BLC bit driver and its associated SDS bit driver located at port 15 in the SDS rack. Was able to establish a loop back at 1615 GMT. Disconnected the 9-pin data cable from the rear of the BLC laptop computer and connected it to another laptop and attempted to do another loop back test. This time a loop back was not successful. This indicated that there was a problem between the BLC laptop computer and the bit driver located at port 15 in the SDS rack. Discovered that the data line from the BLC laptop was plugged into data port 10B. However, the data line which feeds the bit driver at port 15 in the SDS rack was not plugged into port 10B in the SDS patch panel. Instead, the data line was plugged into serial port 23 on the Development system. Disconnected the data cable from serial port 23 on the Development system and plugged it into port 10B on the patch panel at 1645 GMT. Removed the loop back plug and reconnected the BLC bit driver to the BLC RS-232 connector at 1655 GMT. Communications restored between the BLC and its laptop computer at 1700 GMT. RTS.

1/9/98

MW

The BLC screen was not updating. It appeared that no data packets were being received from the BLC. Restarted the BLC computer at 1715 GMT. Still no data packets received after reboot. Power cycled the BLC itself at 1725 GMT. This did not correct the problem. T/S still in progress.

1/5/98

MW

The BLC screen was not updating and no data was being collected. Power cycled the BLC laptop at 1736 GMT. The laptop was not receiving any data from the Ceilometer. Power cycled the BLC itself at 1807 GMT. The laptop resumed communications with the Ceilometer and the display began updating normally.

12/15/97

MW

The BLC computer was locked up this morning. Rebooted the computer at 1430 GMT. RTS at 1435 GMT.

12/11/97

CAM

The laptop display icons and graph were extremely large and could not be interpreted accurately. Exited the program and power cycled the computer at 1446 GMT. RTS status O.K. at 1449 GMT.

12/5/97

CAM

The BLC display was not updating during daily status check at 1630 GMT. The last data packed displayed was at 14:37:30 GMT today. Rebooted computer at 1634 GMT. The display still did not update after reboot with no data packets being received. Power cycled instrument at 1647 GMT. RTS status O.K. at 1651 GMT with display now updating.

11/12/97

CAM

Although the display was updating and cloud heights were indicated in the text box, the graph did not appear normal. Rather than being a continuous graph, the display only showed scattered dots. Exited the clouds program and rebooted computer at 1645 GMT. System then resumed normal operation and graph appeared correct. RTS at 1653 GMT.

11/12/97

CAM

The PC time was approximately 3 minutes fast. PC clock showed 1653 at 1650 GMT. Reset PC clock to 16:50:30 GMT. RTS status O.K.

6/12/97

MW

The BLC clock was reading 5 minutes ahead of WWV time. Used 'alt-T' to enter terminal and used the 'time' command at the prompt to change the time to match WWV at 1447 GMT.

4/23/97

ZW

On Tuesday 4/22/97 at 1730 GMT Conner Flynn and site ops removed the Vaisala Ceilometer from service. Disconnecting power to the Vaisala Ceilometer at the UPS also disconnected power to the Belfort Ceilometer because they both use the same outlet on the UPS. After removing the Vaisala, power to the Belfort was inadvertently left disconnected. Reconnected power to the Belfort at 1338 GMT 4/23/97. Status ok.

3/19/97

ZW

During the daily PM it was found that the Belfort Laser Ceilometer time was 10 min faster than SDS time. Reset the laptop time at 1530 GMT from 0750 GMT to 1530 GMT. Restarted DEMO program but Ceilometer time was still 10 minutes fast. Notified Connor Flynn who supplied a procedure for resetting the Ceilometer BIOS time. Reset the Ceilometer BIOS time at 1908 GMT. Status ok. RTS 1908 GMT.

12/17/96

CAM

FEP data has not updated since 0747 GMT. Rebooted computer at 1415 GMT with no change. T/S found the optical trailer UPS output breaker had tripped. Reset the breaker but it overloaded and tripped again. AERI01 and Ceilometer were on the same breaker. Paired the Ceilometer with WSI and RTS o.k. at 1515 GMT. 800W heater coming on due to cold temperature likely caused the overload.

5/24/96

KT/ZW

PM. Cleaned windows and mirrors from 1330 to 1555 GMT.

2/13/96

GRS

1525 GMT powered down for work on the 60MT. RTS at 1938 GMT.

5/ 1/96

CAM

PM. Powered down to clean mirrors from 1611 to 1634 GMT.

12/27/95

CAM/TMG

1505 GMT powered down for work on the Calibration trailer.

1/29/95

CAM

1430 GMT powered down for work on the Calibration trailer. RTS at 2030 GMT.